# Unveiling the Hidden: Movie Genre and User Bias in Spoiler Detection

### v2

@@ 3.5. Learning and Optimization
4. Experiment
4.1. Experiment Settings
4.1.1. Dataset.
To evaluate our GUSD framework along with 14 other representative baselines on two widely recognized datasets: LCS (Wang et al., 2023) and Kaggle (Wan et al., 2019):
  • Kaggle, introduced in 2019, consists of 573,913 valid reviews, 293,407 users, and 1,572 movies. And about 25.87% (150,924) of the reviews are spoilers.
Note that both datasets include the genre information of all movies, with specific categories defined according to IMDb standards, which facilitates the operation of our GUSD framework. Following MVSD (Wang et al., 2023), we randomly split the reviews into training, validation, and test sets with a ratio of 7:2:1.
4.1.2. Baselines.
To achieve a comprehensive evaluation, we compare GUSD with pre-trained language models, GNN-based models, and task-specific baselines. For the pre-trained language models, the procedure involves feeding the review text into the model, averaging all token embeddings, and then applying two fully connected layers to perform spoiler detection. Regarding the GNN-based models, the graph neural network takes the output of RoBERTa (Liu, 2019) as the initial node features. Below, we provide a concise overview of each baseline method.
  • BERT (Devlin, 2018) is a language model pre-trained on extensive natural language data, using masked language modeling and next sentence prediction tasks.
  • RoBERTa (Liu, 2019) improves upon BERT by eliminating the next sentence prediction task and enhancing masking techniques.
  • BART (Lewis, 2019) is a pre-trained language model that advances traditional autoregressive models through bidirectional encoding and denoising objectives.
  • DeBERTa (He et al., 2021) refines BERT by implementing disentangled attention and an improved mask decoder, making it a more advanced language model.
  • Bge-Large (Xiao et al., 2023) is trained on a comprehensive training dataset C-MTP, combining vast unlabeled data and diverse labeled data.
  • GCN (Kipf and Welling, 2016) is a foundational graph neural network that performs convolutions on graph nodes and their neighbors, effectively propagating information.
  • R-GCN (Schlichtkrull et al., 2018) extends GCN to handle multi-relational graphs by incorporating relation-specific weights.
  • GAT (Veličković et al., 2017) is a graph neural network that applies attention mechanisms to dynamically assign importance to neighboring nodes.
  • SimpleHGN (Lv et al., 2021) is tailored for heterogeneous graphs, integrating multiple types of nodes and edges with a shared embedding space and adaptive aggregation strategies.
  • GPS (Rampášek et al., 2022) propose a recipe to build a general, powerful, scalable graph Transformer with linear complexity.
  • HGT (Hu et al., 2020) design node- and edge-type dependent parameters to characterize the heterogeneous attention over each edge for modeling Web-scale heterogeneous graphs.
  • DNSD (Chang et al., 2018) is a spoiler detection method that employs a CNN-based genre-aware attention mechanism.
  • SpoilerNet (Wan et al., 2019) uses a hierarchical attention network and GRU alongside item and user bias terms for spoiler detection.
  • MVSD (Wang et al., 2023) leverages external movie knowledge and user networks to detect spoilers.
4.2. Main Results
@@
  • GUSD consistently outperforms all baselines across both datasets. Specifically, compared with the previous state-of-the-art method MVSD (Wang et al., 2023), GUSD achieves 6.1% higher Binary-F1, 5.5% higher AUC, and 2.0% higher accuracy on the LCS dataset, as well as 15.2% higher Binary-F1, 11.6% higher AUC, and 6.1% higher accuracy on the Kaggle dataset. These improvements are statistically significant.
@@
  • Compared with DNSD (Chang et al., 2018), which also focuses on genre features of the reviews, GUSD surpasses DNSD across all three metrics in both datasets, further proving the effectiveness and robustness of our global-aware genreformer and GMoE methods.
  • Both SpoilerNet (Wan et al., 2019) and GUSD utilize user bias, but GUSD outperforms SpoilerNet in all three metrics across both datasets. This demonstrates that our dynamic graph pretraining can better identify the latent behavior pattern of whether a user is likely to post spoilers.
4.3. Ablation Study
@@
  • For the RetGAT component, we evaluated two variations: using an approximate method to compute $k$-hop neighbors (Atwood and Towsley, 2016; Wang et al., 2020) and replacing our RetGAT with a standard GAT. Both variations lead to a drop in performance, indicating the necessity of our RetGAT design for capturing appropriate graph structures.
  • To investigate the impact of GMoE, we replace it with a simple MLP, a traditional MoE (Shazeer et al., 2017), and a Soft-MoE (Puigcerver et al., 2023). Note that we set the number of experts of the traditional MoE and Soft-MoE to be the same as in GMoE, i.e., the number of genres. From the results shown in Table 2, the full model with GMoE achieves the best performance, highlighting the effectiveness of our GMoE design in handling explicit genre information. Moreover, replacing the GMoE with an MLP performs the worst, proving the rationality of using genre information.
@@
Table 2. Ablation study of GUSD on Kaggle Dataset. Bold indicates the best performance, underline the second best.
@@
4.4. GMoE Study
In this section, we conduct further experiments on the GMoE to better understand its effectiveness compared to other previous MoE methods such as traditional MoE (Shazeer et al., 2017) and Soft-MoE (Puigcerver et al., 2023).
4.4.1. Different numbers of Experts.
@@ 4.4. GMoE Study
Figure 4. Performance comparison of different numbers of experts in traditional MoE and Soft-MoE. Note that 21 is the number of genres. The results indicate that GMoE outperforms other variants irrespective of the number of experts.
Figure 5. T-SNE visualization of the features processed by GMoE. Different colors represent different genres, indicating distinct clustering of features according to genres.
Table 3. Evaluation of user bias effectiveness. The user bias pre-trained on Dynamic Graph achieves significantly higher accuracy in predicting user labels based on spoiler review percentage compared to raw user embeddings.
@@
4.4.2. T-SNE Visualization of Features.
Table 4. Examples of the performance of top-2 baselines and GUSD. Underlined parts indicate the plots. ”Key Information” indicates the most helpful information from other sources when detecting spoilers.
@@
Figure 6. When randomly removing edges from the graph, randomly setting elements of graph features to zero, and reducing the number of labeled samples in the training set, the performance of GUSD gradually declines with these incremental ablations. However, the fact that the model maintains a reasonable level of performance despite these perturbations indicates the robustness of our method.
4.5. User Bias Analysis
@@
4.6. Robustness Study
In the realm of social media analysis, particularly for tasks like spoiler detection, obtaining pure and clean data is a significant challenge. Spoiler detection often deals with information that is noisy and heterogeneous. As highlighted by previous studies on bot detection (Cai et al., 2024; Liu et al., 2023), it is crucial to investigate the robustness of models under conditions where information is incomplete or noisy.
@@ 4.7. Case Study
5. Ethical Considerations
We propose GUSD as an initial filtering tool rather than a final arbiter. Despite attaining state-of-the-art performance, GUSD remains imperfect and must be employed judiciously, in conjunction with human moderators to oversee or halt questionable movie reviews. Additionally, GUSD may reflect the biases embedded in its underlying datasets and models. For example, pretrained language models might perpetuate unwanted social biases and stereotypes (Li et al., 2022; Nadeem et al., 2020). It is within the scope of future research to integrate bias detection and mitigation strategies from machine learning into spoiler detection systems. Furthermore, due to the nature of the task, the dataset includes potentially offensive language, which warrants careful consideration.
6. Conclusion
Appendix A Implementation Details
We use PyTorch (Paszke et al., 2019), scikit-learn (Pedregosa et al., 2011), PyTorch Geometric (Fey and Lenssen, 2019), MoE (Rau, 2019), soft-mixture-of-experts (Puigcerver et al., 2023) and Transformers  (Wolf et al., 2020) library to implement our proposed GUSD. We conduct our experiments on a cluster with 4 Tesla V100 GPUs with 32 GB memory, 16 CPU cores, and 377GB CPU memory. Code, data, hyperparameters and trained models will be made publicly available.
Appendix B Hyperparameters
@@
Appendix C k-hop Neighbors Algorithm
Previous studies (Atwood and Towsley, 2016; Wang et al., 2020) have used the $k$-th power of the adjacency matrix to capture $k$-hop visibility. This method does not strictly define $k$-hop neighbors as nodes reachable in exactly k steps. To address this, we introduce a new GPU-friendly $k$-hop Algorithm that quickly and accurately computes $k$-hop neighbors using adjacency matrices. The details are illustrated in Algorithm 1.
@@
Table 5. Hyperparameter settings of GUSD in Kaggle
Hyperparameter Kaggle LCS
optimizer AdamW AdamW
learning rate 1e-4 1e-4
weight decay 1e-5 1e-5
CrossEntropyLoss weight 1.3 1.3
epochs 20 20
dropout 0.3 0.3
batch size 128 128
early stop patience 3 3
LM Bge-Large (Xiao et al., 2023) Roberta (Liu, 2019)
GNN layer 2 3
GNN input dim 1024 768
GNN hidden dim 1024 768
GNN output dim 1024 768
RetGAT hop num 2 2
RetGAT hop decay $\alpha$ 0.3 0.3
meta dim 3 6
meta encoder hidden dim 1024 768
meta encoder output dim 1024 768
TRM heads 4 4
TRM layers 2 2
GMoE layer 2 2
GMoE hidden dim 1024 768
GMoE output dim 1024 768
GMoE experts number 21 28

### v3

@@ 4.1 Experiment Settings
4.2 Main Results
@@
  • GUSD consistently outperforms all baselines across both datasets. Specifically, compared with the previous state-of-the-art method MVSD [38], GUSD achieves 6.1% higher Binary-F1, 5.5% higher AUC, and 2.0% higher accuracy on the LCS dataset, as well as 15.2% higher Binary-F1, 11.6% higher AUC, and 6.1% higher accuracy on the Kaggle dataset. These improvements are statistically significant.
@@
  • Compared with DNSD [6], which also focuses on genre features of the reviews, GUSD surpasses DNSD across all three metrics in both datasets, further proving the effectiveness and robustness of our global-aware genreformer and GMoE methods.
  • Both SpoilerNet [36] and GUSD utilize user bias, but GUSD outperforms SpoilerNet in all three metrics across both datasets. This demonstrates that our dynamic graph pretraining can better identify the latent behavior pattern of whether a user is likely to post spoilers.
4.3 Ablation Study
@@
  • For the RetGAT component, we evaluated two variations: using an approximate method to compute $k$-hop neighbors [1, 37] and replacing our RetGAT with a standard GAT. Both variations lead to a drop in performance, indicating the necessity of our RetGAT design for capturing appropriate graph structures.
  • To investigate the impact of GMoE, we replace it with a simple MLP, a traditional MoE [33], and a Soft-MoE [29]. Note that we set the number of experts of the traditional MoE and Soft-MoE to be the same as in GMoE, i.e., the number of genres. From the results shown in Table 2, the full model with GMoE achieves the best performance, highlighting the effectiveness of our GMoE design in handling explicit genre information. Moreover, replacing the GMoE with an MLP performs the worst, proving the rationality of using genre information.
@@
Table 2: Ablation study of GUSD on Kaggle Dataset. Bold indicates the best performance, underline the second best.
@@
4.4 GMoE Study
In this section, we conduct further experiments on the GMoE to better understand its effectiveness compared to other previous MoE methods such as traditional MoE [33] and Soft-MoE [29].
4.4.1 Different numbers of Experts.
@@ 4.4 GMoE Study
Figure 4: Performance comparison of different numbers of experts in traditional MoE and Soft-MoE. Note that 21 is the number of genres. The results indicate that GMoE outperforms other variants irrespective of the number of experts.
Figure 5: T-SNE visualization of the features processed by GMoE. Different colors represent different genres, indicating distinct clustering of features according to genres.
4.4.2 T-SNE Visualization of Features.
5 Conclusion
{credits}
5.0.1 Acknowledgements
This work was supported by the National Nature Science Foundation of China (No. 62272374, No. 62192781 ), the Natural Science Foundation of Shaanxi Province (2024JC-JCQN-62), the National Nature Science Foundation of China (No. 62202367, No. 62250009), the Key Research and Development Project in Shaanxi Province No. 2023GXLH-024, Project of China Knowledge Center for Engineering Science and Technology, and Project of Chinese academy of engineering “The Online and Offline Mixed Educational Service System for ‘The Belt and Road’ Training in MOOC China”, and the K. C. Wong Education Foundation. Lastly, we would like to thank all LUD lab members for fostering a collaborative research environment.
Appendix 0.A Implementation Details
We use PyTorch [27], scikit-learn [28], PyTorch Geometric [11], MoE [31], soft-mixture-of-experts [29] and Transformers  [39] library to implement our proposed GUSD. We conduct our experiments on a cluster with 4 Tesla V100 GPUs with 32 GB memory, 16 CPU cores, and 377GB CPU memory. Code, data, hyperparameters and trained models will be made publicly available.
Appendix 0.B More Experiments
Table 3: Evaluation of user bias effectiveness. The user bias pre-trained on Dynamic Graph achieves significantly higher accuracy in predicting user labels based on spoiler review percentage compared to raw user embeddings.
@@ Appendix 0.B More Experiments
Table 4: Examples of the performance of top-2 baselines and GUSD. Underlined parts indicate the plots. "Key Information" indicates the most helpful information from other sources when detecting spoilers.
@@
Figure 6: When randomly removing edges from the graph, randomly setting elements of graph features to zero, and reducing the number of labeled samples in the training set, the performance of GUSD gradually declines with these incremental ablations. However, the fact that the model maintains a reasonable level of performance despite these perturbations indicates the robustness of our method.
0.B.1 User Bias Analysis
@@
0.B.2 Robustness Study
In the realm of social media analysis, particularly for tasks like spoiler detection, obtaining pure and clean data is a significant challenge. Spoiler detection often deals with information that is noisy and heterogeneous. As highlighted by previous studies on bot detection [4, 20], it is crucial to investigate the robustness of models under conditions where information is incomplete or noisy.
@@ 0.B.3 Case Study
Table 5: Hyperparameter settings of GUSD in Kaggle
Hyperparameter Kaggle LCS
optimizer AdamW AdamW
learning rate 1e-4 1e-4
weight decay 1e-5 1e-5
CrossEntropyLoss weight 1.3 1.3
epochs 20 20
dropout 0.3 0.3
batch size 128 128
early stop patience 3 3
LM Bge-Large [41] Roberta [19]
GNN layer 2 3
GNN input dim 1024 768
GNN hidden dim 1024 768
GNN output dim 1024 768
RetGAT hop num 2 2
RetGAT hop decay $\alpha$ 0.3 0.3
meta dim 3 6
meta encoder hidden dim 1024 768
meta encoder output dim 1024 768
TRM heads 4 4
TRM layers 2 2
GMoE layer 2 2
GMoE hidden dim 1024 768
GMoE output dim 1024 768
GMoE experts number 21 28
Appendix 0.C Hyperparameters
@@
Appendix 0.D k-hop Neighbors Algorithm
Previous studies [1, 37] have used the $k$-th power of the adjacency matrix to capture $k$-hop visibility. This method does not strictly define $k$-hop neighbors as nodes reachable in exactly k steps. To address this, we introduce a new GPU-friendly $k$-hop Algorithm that quickly and accurately computes $k$-hop neighbors using adjacency matrices. The details are illustrated in Algorithm 1.
Appendix 0.E Metadata Collected for Nodes in Datasets
Here we provide the details of metadata collected for nodes in two datasets in Table 6
Table 6: Comparison of metadata fields used in the LCS and IMDB datasets.
Entity Field Type LCS IMDB
Movie year int
isAdult bool
runtime int
rating float
numVotes int
genres list
User badge_count int
review_count int
avg_helpful float
avg_total float
avg_score float
avg_rating float
avg_length float
Review helpful_votes int
total_votes int
score float
rating float
review_text_length int